OpenStudy (anonymous):

Write the number 117 as a product of prime numbers

OpenStudy (unklerhaukus):

there are only two

OpenStudy (anonymous):

what are they?

OpenStudy (unklerhaukus):

try dividing 117 by some various smallish prime numbers , if your answer has no remainder then that number is a factor 117/2=58 remainder 1 so two is not a factor, ....,

OpenStudy (anonymous):

@preston_hunter

OpenStudy (anonymous):

u got it .........

OpenStudy (anonymous):

the sum of the digits of 117 is divisible by 9 and 117/9=13 so we have 117=3x3x13 and the canonical form is 117=3^2x13